# Universal bounds on optimization of free energy harvesting

### 20 Dec 2023 v3

@@ Abstract
To explore the performance of bacteriorhodopsin under different conditions, we vary the membrane electrical potential $\Delta\psi$ between $-75$ and $350$ mV, while using a realistic fixed $\Delta\text{pH}=-0.6$ [38]. We show $\dot{\mathcal{G}}^{\mathrm{tot}}$, the actual harvesting rate at different potentials, as a black line in Fig. 1 (units $k_{B}T/$sec). At a plausible in vivo $\Delta\psi=120$ mV [38], the model exhibits a steady-state current of 11.5 protons/sec, each proton carrying 6.1 $k_{B}T$ of free energy, corresponding to ${\dot{\mathcal{G}}^{\mathrm{tot}}}\approx 70\,k_{B}T$/sec. The largest output is achieved near the in vivo potential: at lower potentials, the cycle current saturates while the free energy per proton drops, and at higher potentials the pump stalls.
@@
Fig. 1 shows $\mathscr{G}$, the maximum $\dot{\mathcal{G}}^{\mathrm{tot}}$ achievable by optimizing each reversible transition. For each transition, we also show the efficiency $\dot{\mathcal{G}}^{\mathrm{tot}}/\mathscr{G}\leq 1$, the ratio of the actual and maximum harvesting rate.
@@
Several transitions, such as $K\leftrightarrow L$,$L\leftrightarrow M_{1}$,$M_{1}\leftrightarrow M_{2}$, are remarkably efficient ($\geq 85\%$) near in vivo membrane potentials. The reprotonation step $N\leftrightarrow O$ is the least efficient ($\sim 40\%$) and also has the slowest kinetics of the 5 transitions shown in Fig. 1. This suggests that $N\leftrightarrow O$ is a bottleneck whose optimization can have a big impact on the harvesting rate, while optimization of other (non-bottleneck) transitions has a more limited effect.
@@
Observe that $\mathscr{G}$ for $M_{1}\leftrightarrow M_{2}$ does not depend on $\Delta\psi$. This is because $\mathscr{G}$ is a function of baseline properties, which do not depend on the membrane potential when $M_{1}\leftrightarrow M_{2}$ is treated as control. Conversely, $M_{1}\leftrightarrow M_{2}$ as control transition can be optimized by varying the membrane potential and by simultaneously scaling up the forward/backward kinetics. Our results show that this transition is very close to optimal at in vivo membrane potentials and kinetic timescales.
@@
Optimal distributions, $\bm{p}^{*}$, are also obtained (example in Fig. 1, lower left). We find a consistent shift toward state $O$, which accelerates the reset of the cycle and increases the flux across the photon-absorbing transition $O\to K$.
In SM2 [31], we study bacteriorhodopsin under further kinetic and thermodynamic constraints. We also consider a complementary analysis that uses (6), in which we treat the full cycle as baseline and allow for arbitrary additional control processes.
@@
Limiting regimes.— Our results are stated via an optimization problem that generally does not have a closed-form solution. In our second set of results, we identify closed-form expressions in three physically meaningful regimes. For simplicity, here we focus on the simplified objective (6). (See SM3 [31] for derivations, including validity conditions for the three regimes.)
@@
Linear response (LR) applies when the optimal distribution $\bm{p}^{*}$ is close to steady-state distribution of the baseline rate matrix $R^{b}$. Suppose that $R^{b}$ is irreducible and has a unique steady state ${\bm{\pi}^{b}}$ with full support. We introduce the “additive reversibilization” of $R^{b}$, $A_{ij}=(R^{b}_{ij}+R^{b}_{ji}\pi^{b}_{i}/\pi^{b}_{j})/2$, which obeys detailed balance (DB) for the steady-state distribution ${\bm{\pi}^{b}}$ and has the same dynamical activity [46] on all edges as $R^{b}$. $A$ is a DB version of $R^{b}$, being equal to $R^{b}$ when the latter obeys DB [47, 48]. Let $\boldsymbol{u}^{\alpha}$ indicate the $\alpha$-th right eigenvector of $A$ normalized as $\sum_{i}({u^{a}_{i}})^{2}/\pi^{b}_{i}=1$, and $\lambda_{\alpha}$ the corresponding real-valued eigenvalue ($\lambda_{1}=0$).
The LR solution for maximum harvesting rate and optimal distribution is
$\mathscr{G}\approx\dot{\mathcal{G}}^{b}({\bm{\pi}^{b}})+\beta\sum_{\alpha>1}\frac{\left|\Omega_{\alpha}\right|^{2}}{-\lambda_{\alpha}},\;\;\bm{p}^{*}\approx{\bm{\pi}^{b}}+\beta\sum_{\alpha>1}\frac{\Omega_{\alpha}}{-\lambda_{\alpha}}\boldsymbol{u}^{\alpha}$ (9)
@@
The Deterministic (D) regime applies when the nonlinear information-theoretic term in (8) is much smaller than the linear thermodynamic term. We can then ignore the former, turning (8) into a simple linear optimization. This gives the approximate solution $\mathscr{G}\approx\phi_{{i}^{*}}$, ${p}^{*}_{i}\approx\delta_{{i}^{*}i}$, where ${i}^{*}=\arg\max_{i}\phi_{i}$ is the optimal mesostate. This solution concentrates probability on a single mesostate, effectively ignoring the cost of maintaining this low-entropy distribution.
The Near-Deterministic (ND) regime lies between Linear Response and Deterministic ones. By perturbing $\bm{p}^{*}$ around $\delta_{{i}^{*}i}$, we can decouple the values of $p_{i}$ in the objective function (6). This gives the approximate solution $\mathscr{G}\approx\phi_{{i}^{*}}+{\beta}^{-1}\sum_{i\neq{i}^{*}}R^{b}_{i{i}^{*}}(\ln{p}^{*}_{i}-1)$, ${p}^{*}_{i(\neq{i}^{*})}\approx{R^{b}_{i{i}^{*}}}/({\beta(\phi_{{i}^{*}}-\phi_{i})+R^{b}_{{i}^{*}{i}^{*}}})$, ${p}^{*}_{{i}^{*}}=1-\sum_{i:i\neq{i}^{*}}{{p}^{*}_{i}}$.
@@
Figure 2: Left: unicyclic system where free energy $\mathscr{G}$ is harvested by a single transition. Middle: Log-log plot of $\mathscr{G}$ and its approximations for different $\Theta$ ($n=5$, $\beta=1$). Right: exact and approximate optimal solution $\bm{p}^{*}$ using ND (top) and LR (bottom) approximations. The optimal state 1 is in the middle of the histogram.
@@
Example: unicyclic network.— We illustrate our closed-form solutions using a minimal model of a unicyclic system with $n$ states (Fig. 2). The baseline dynamics involve diffusion across a 1-dimensional ring, with left and right jump rates set to 1. The baseline steady state is uniform with no cyclic current.
Suppose $\Theta$ of free energy is harvest during the jump $1\to 2$ (${g}^{b}_{21}=-{g}^{b}_{12}=\Theta$), so that harvesting requires a cyclic current. This system can be interpreted as an idealized model of a biomolecular harvesting cycle, such as a transporter.
In SM4 [31], we study the maximum harvesting rate $\mathscr{G}$ and the optimal distribution in the LR, ND, and D regimes. Here we summarize the results (Fig. 2). For small $\Theta$, LR applies and $\mathscr{G}\approx\beta\Theta^{2}{(n-1)}/{4n^{2}}$. For large $\Theta$, the D regime is relevant and the optimal distribution concentrates on the optimal state ${i}^{*}=1$ and $\mathscr{G}\approx\Theta$. At intermediate $\Theta$, the ND regime gives $\mathscr{G}\approx\Theta-{\beta}^{-1}\big{\{}2+\ln[2(\beta\Theta-1)(\beta\Theta-2)]\big{\}}$.
@@
Interestingly, in the LR regime at small $\Theta$, $\mathscr{G}$ scales quadratically in $\Theta$ and essentially linearly in $1/n$ and $\beta$. This scaling arises because the optimal strategy balances harvesting ($\mathscr{G}$ contribution) with the thermodynamic cost of maintaining a low entropy distribution ($\beta$ and $1/n$ contributions). For larger $\Theta$ (ND and D regimes), $\mathscr{G}$ scales linearly in $\Theta$ and loses its linear dependence on $\beta$ and $1/n$. Thus, in the presence of large free energy sources, the thermodynamic cost of maintaining a low entropy distribution has a minor effect on the optimal strategy.
In SM4 [31], we also discuss a different unicyclic model, where harvesting depends on internal fluxes within a single mesostate ($\dot{g}^{b}_{1}>0$).

### 12 Mar 2024 v4

@@ IV Bacteriorhodopsin
To explore the performance of bacteriorhodopsin under different conditions, we vary the membrane electrical potential $\Delta\psi$ between $-75$ and $350$ mV, while using a realistic fixed $\Delta\text{pH}=-0.6$ [34]. We show the actual harvesting rate ($\dot{\mathcal{G}}^{\mathrm{tot}}$ in units of $k_{B}T/$sec) at different potentials as a black line in Fig. 2 (a). At a plausible in vivo $\Delta\psi=120$ mV [34], the model exhibits a steady-state current of 11.5 protons/sec, each proton carrying 6.1 $k_{B}T$ of free energy, corresponding to ${\dot{\mathcal{G}}^{\mathrm{tot}}}\approx 70\,k_{B}T$/sec. The largest output is achieved near the in vivo potential: at lower potentials, the cycle current saturates while the free energy per proton drops, and at higher potentials the pump stalls.
@@
Figure 2: (a) Comparison of the actual harvesting rate $\dot{\mathcal{G}}^{\mathrm{tot}}$ at different electrical potentials $\Delta\psi$, versus maximum rate $\mathscr{G}$ achieved by optimizing five intermediate transitions (color scheme from Fig. 1 Right). (b) Efficiency $\dot{\mathcal{G}}^{\mathrm{tot}}/\mathscr{G}$ computed while separately optimizing each transition, with colors as in (a). (c) The actual steady state $\bm{\pi}$ versus the optimal distribution $\bm{p}^{*}$ when optimizing the $N\leftrightarrow O$ transition (at $\Delta\psi=120$ mV).
@@
Fig. 2 (a) shows $\mathscr{G}$, the maximum $\dot{\mathcal{G}}^{\mathrm{tot}}$ achievable by optimizing each reversible transition. In Fig. 2 (b), we also show the efficiency $\dot{\mathcal{G}}^{\mathrm{tot}}/\mathscr{G}\leq 1$ for each transition, that is the ratio of the actual and maximum harvesting rate.
@@
Several transitions, such as $K\leftrightarrow L$,$L\leftrightarrow M_{1}$,$M_{1}\leftrightarrow M_{2}$, are remarkably efficient ($\geq 85\%$) near in vivo membrane potentials. The reprotonation step $N\leftrightarrow O$ is the least efficient ($\sim 40\%$) and also has the slowest kinetics of the 5 transitions studied in Fig. 2. This suggests that $N\leftrightarrow O$ is a bottleneck whose optimization can have a big impact on the harvesting rate, while optimization of other non-bottleneck transitions has a more limited effect.
@@
Observe that $\mathscr{G}$ for $M_{1}\leftrightarrow M_{2}$ does not depend on $\Delta\psi$. This is because $\mathscr{G}$ is a function of baseline properties, which do not depend on the membrane potential when $M_{1}\leftrightarrow M_{2}$ is treated as control. Conversely, $M_{1}\leftrightarrow M_{2}$ as control transition can be optimized by varying the membrane potential and/or scaling up the forward/backward rates. Our results show that this transition is very close to optimal at in vivo membrane potentials and kinetic timescales.
@@
Optimal distributions $\bm{p}^{*}$ are also obtained, with a typical one shown in Fig. 2 (c). We find a consistent shift toward state $O$, which accelerates the reset of the cycle and increases the flux across the photon-absorbing transition $O\to K$.
In SM-II [27], we illustrate how the efficiency of bacteriorhodopsin transitions can be evaluated under other types of constraints, including constraints on thermodynamic affinity, dynamical activity, and kinetics.
Figure 3: (a) Comparison of the actual harvesting rate $\dot{\mathcal{G}}^{\mathrm{tot}}$ at different electrical potentials $\Delta\psi$, versus maximum rate $\mathscr{G}$ achieved by fixing the bacteriorhodopsin cycle as baseline and allowing any additional transitions as control. (b) Efficiency of the actual bacteriorhodopsin cycle with respect to the optimized cycle. (c) The actual steady state $\bm{\pi}$ and optimal distribution $\bm{p}^{*}$ (at $\Delta\psi=120$ mV).
As a final analysis, instead of optimizing individual existing transitions in the bacteriorhodopsin cycle, we ask to what extent the harvesting rate can be increased by any additional control processes. For example, this could involve an additional enzyme that shifts the cycle’s steady state by permitting a new transition between distant states (e.g. $L\leftrightarrow N$), possibly yielding an increase of the proton pumping rate.
In this case, we treat the entire bacteriorhodopsin system as the baseline, and we do not introduce any additional constraints on the steady-state distribution or the control fluxes. Then, as shown in Sec. SM-I.4 [27], the objective in (10) is achieved in the limit of fast control, and the maximum harvesting rate can be found by solving the simplified optimization problem (11).
For this setup, Fig. 3 (a) shows the baseline (actual) harvesting rate $\dot{\mathcal{G}}^{\mathrm{tot}}$ and the maximum harvesting rate $\mathscr{G}$ at varying $\Delta\psi$. Interestingly, both peak at around the in vivo values of the membrane potential. In Fig. 3 (b), we show that the actual bacteriorhodopsin cycle harvests approximately 50% of the fundamental bound given by $\mathscr{G}$ (at in vivo values of $\Delta\psi$). This suggests that bacteriorhodopsin is remarkably close to optimal, relative to improvements that could be achieved by introducing any additional control processes.
We also show the actual steady-state distribution and the optimal distribution $\bm{p}^{*}$ in Fig. 3 (c). The optimal distribution increases the probability of state $O$, similar to the optimal distribution found by optimizing the $N\leftrightarrow O$ transition, shown in Fig. 2 (c). However, unlike Fig. 2 (c), where most of the extra probability is taken from state $N$, in Fig. 3 (c) the probability is drawn more uniformly from other states in the cycle, indicating the presence of distributed control.
V Limiting regimes
@@
Our results are stated via an optimization problem that generally does not have a closed-form solution. In our second set of results, we identify closed-form expressions in three physically meaningful regimes. For simplicity, here we focus on the simplified objective (11). See SM-III [27] for detailed derivations, including analysis of the conditions under which each of these three approximation are be valid.
@@
Linear response (LR) applies when the optimal distribution $\bm{p}^{*}$ is close to the steady-state distribution of the baseline rate matrix $R^{b}$. Suppose that $R^{b}$ is irreducible and has a unique steady state ${\bm{\pi}^{b}}$ with full support. We introduce the “additive reversibilization” of $R^{b}$,
$A_{ij}=(R^{b}_{ij}+R^{b}_{ji}\pi^{b}_{i}/\pi^{b}_{j})/2\,.$ (14)
The rate matrix $A$ obeys detailed balance (DB) for the steady-state distribution ${\bm{\pi}^{b}}$ and has the same dynamical activity [42] on all edges as $R^{b}$. $A$ may be considered as a DB version of $R^{b}$, and it is equal to $R^{b}$ when the latter obeys DB [43, 44]. Let $\boldsymbol{u}^{\alpha}$ indicate the $\alpha$-th right eigenvector of $A$ normalized as $\sum_{i}({u^{\alpha}_{i}})^{2}/\pi^{b}_{i}=1$, and $\lambda_{\alpha}$ the corresponding real-valued eigenvalue ($\lambda_{1}=0$). The LR solution for the maximum harvesting rate and the optimal distribution is
[…]
@@
The Deterministic (D) regime applies when the nonlinear information-theoretic term in (13) is much smaller than the linear thermodynamic term. We can then ignore the former, turning (13) into a simple linear optimization. This gives the approximate solution
$\mathscr{G}\approx\phi_{{i}^{*}}\quad\qquad{p}^{*}_{i}\approx\delta_{{i}^{*}i}\,$ (16)
where ${i}^{*}=\arg\max_{i}\phi_{i}$ is the optimal mesostate. This solution concentrates probability on a single mesostate, effectively ignoring the cost of maintaining this low-entropy distribution.
The Near-Deterministic (ND) regime lies between Linear Response and Deterministic ones. By perturbing $\bm{p}^{*}$ around $\delta_{{i}^{*}i}$, we can decouple the values of $p_{i}$ in the objective function (11). The maximal harvesting rate and optimal distribution in this regime are then given by
[…]
@@
Figure 4: (a) Unicyclic system where free energy $\Theta$ is harvested by a single transition. (b) Unicyclic system where free energy per unit time $\theta$ is harvested when the system is in a particular mesostate.
VI Example: unicyclic systems
@@
We illustrate our closed-form solutions using two simple models, both based on a unicyclic system with $n$ states. The baseline dynamics involve diffusion across a 1-dimensional ring, with left and right jump rates set to 1. The baseline steady state is a uniform distribution, $\pi^{b}_{i}=1/n$, with no cyclic current. We assume a uniform free energy function, $f_{i}=0$ for all $i$.
We consider two different scenarios. In the first scenario, shown schematically in Fig. 4 (a), $\Theta$ of free energy is harvested each time the system carries out the transition $1\to 2$, so
${g}^{b}_{21}=-{g}^{b}_{12}=\Theta\,,$
and ${g}^{b}_{ij}=\dot{g}^{b}_{i}=0$ otherwise. This scenario may be interpreted as an idealized model of a biomolecular harvesting cycle, such as a transporter. In the second scenario, shown schematically in Fig. 4 (b), free energy is harvested at a rate of $\theta$ per unit time when the system is located in one particular mesostate ${i}^{*}=1$, so
$\dot{g}^{b}_{1}=\theta\,,$ (18)
and ${g}^{b}_{ij}=\dot{g}^{b}_{i}=0$ otherwise. This scenario may be interpreted as an idealized model of error correction or self-assembly, where free energy can only be harvested when the system is in some particular functional mesostate.
For both scenarios, we evaluate the maximum harvesting rate $\mathscr{G}$ and the optimal distributions achievable by adding any possible control to the system, without constraints. We report exact values found by numerical optimization of Eq. (13), as well as the LR, ND, and D approximations described above. To calculate the LR values, we exploit the fact that the baseline unicyclic rate matrix is a circulant matrix with a simple eigendecomposition [45]. Full details of the derivations for the two scenarios are provided in SM-IV [27] and SM-IV.3 [27], respectively.
Figure 5: (a) Maximum harvesting rate $\mathscr{G}$ for the unicyclic system from Fig. 4 (a), as a function of supplied free energy $\Theta$. Exact value is found numerically, LR, D, and ND are calculated using approximations described in the text. Exact and approximate optimal distributions $\bm{p}^{*}$ in ND (b) and LR (c) regimes are shown, with the optimal state ${i}^{*}=1$ located in the middle of the histograms. (d) $\mathscr{G}$ and its LR approximation for different $\Theta$ and system sizes $n$.
Figure 6: (a) Maximum harvesting rate $\mathscr{G}$ for the unicyclic system from Fig. 4 (b), as a function of supplied free energy rate $\theta$. Exact value is found numerically, LR, D, and ND are calculated using approximations described in the text. Exact and approximate optimal distributions $\bm{p}^{*}$ in ND (b) and LR (c) regimes are shown, with the optimal state ${i}^{*}=1$ located in the middle of the histograms. (d) $\mathscr{G}$ and its LR approximation for different $\theta$ and system sizes $n$.
We first report results for the first scenario from Fig. 4 (a), where free energy is harvested during the transition $2\to 1$. Observe that baseline harvesting rate vanishes, $\dot{\mathcal{G}}^{b}({\bm{\pi}^{b}})=0$, since harvesting free energy requires a cyclic current. In Fig. 5 (a), we plot the maximum harvesting rate $\mathscr{G}$ and its approximations as a function of the supplied free energy $\Theta$. For small $\Theta$, LR applies and the maximum harvesting rate is
$\mathscr{G}\approx\beta\Theta^{2}{(n-1)}/{4n^{2}}.$ (19)
The optimal distribution in the LR regime, shown in Fig. 5 (c), builds up in equal increments starting from $i={i}^{*}+1$ until the optimal state ${i}^{*}=1$, after which it drops sharply. For large $\Theta$, the D regime is relevant and the optimal distribution concentrates on the optimal state ${i}^{*}=1$, so
$\mathscr{G}\approx\Theta\,.$ (20)
At intermediate $\Theta$, the ND regime applies, which gives
$\mathscr{G}\approx\Theta-{\beta}^{-1}\big{\{}2+\ln[2(\beta\Theta-1)(\beta\Theta-2)]\big{\}}.$ (21)
The optimal distribution in the ND regime, shown in Fig. 5 (b), allocates ${p}^{*}_{{i}^{*}-1}=1/(\beta\Theta-2)$, ${p}^{*}_{{i}^{*}+1}=1/(2\beta\Theta-2)$ and the rest to the optimal state ${p}^{*}_{{i}^{*}}$.
Next, we consider the second scenario from Fig. 4 (b), where free energy is harvested when the system is in the optimal mesostate ${i}^{*}=1$. Observe that the uniform baseline steady state assigns $1/n$ probability to the optimal state, thus in this scenario the baseline harvesting rate is $\dot{\mathcal{G}}^{b}({\bm{\pi}^{b}})=\theta/n$. To facilitate comparison with the first scenario, we focus on the increase of the maximum harvesting rate relative to baseline,
$\Delta\mathscr{G}:=\mathscr{G}-\dot{\mathcal{G}}^{b}({\bm{\pi}^{b}})=\mathscr{G}-\theta/n.$ (22)
In Fig. 6 (a), we show $\Delta\mathscr{G}$ and its approximations as a function of the free energy supply rate $\theta$. For small $\theta$, LR applies and the maximum harvesting rate is
$\Delta\mathscr{G}\approx\beta\theta^{2}/48\,.$ (23)
The optimal distribution in the LR regime, shown in Fig. 6 (c), is symmetric about the optimal state ${i}^{*}=1$. For large $\theta$, the D regime is relevant and the optimal distribution concentrates on the optimal state ${i}^{*}=1$, so
$\Delta\mathscr{G}\approx\theta-\theta/n\,.$ (24)
At intermediate $\theta$, the ND regime applies, giving
$\Delta\mathscr{G}\approx\theta-\theta/n-2{\beta}^{-1}\left[1+\ln\left(\beta\theta-2\right)\right].$ (25)
The optimal distribution in the ND regime, shown in Fig. 6 (b), allocates ${p}^{*}_{{i}^{*}-1}={p}^{*}_{{i}^{*}+1}=1/(\beta\theta-2)$ and the rest to ${p}^{*}_{{i}^{*}}$.
@@
There are some similarities among the two harvesting scenarios. For both scenarios, in the LR regime, the increase of the harvesting rate scales quadratically in the supplied free energy ($\Theta$ or $\theta$) and linearly in inverse temperature $\beta$. This scaling reflects the fact that the optimal strategy has to balance harvesting ($\Theta$ or $\theta$ contributions) with the thermodynamic cost of maintaining a low entropy $\bm{p}^{*}$ ($\beta$ contributions). In the ND and D regimes, $\mathscr{G}$ scales linearly in the supplied free energy and loses its linear dependence on $\beta$. Thus, outside of LR, the thermodynamic cost of maintaining a low entropy distribution has a minor effect on the optimal strategy.
There are also important differences between the two scenarios. For the first scenario, the optimal strategy maintains an asymmetric $\bm{p}^{*}$, thereby generating a net flux across the transition $2\to 1$. In the LR regime, the cost of maintaining this asymmetric distribution grows with the system size $n$, therefore the maximum harvesting rate in Eq. (19) scales as $\sim O(n^{-1})$. This is shown in Fig. 5 (d), where we plot $\mathscr{G}$ and its LR approximation at various $\Theta$ and $n$. For the second scenario, the optimal strategy maintains a peaked but symmetric $\bm{p}^{*}$. Remarkably, the cost of maintaining this distribution does not depend on system size $n$. This is shown in Fig. 6 (d), where we plot $\Delta\mathscr{G}$ at various $\theta$ and $n$.
VII Discussion
In this paper, we consider the problem of optimizing free energy harvesting in a nonequilibrium steady-state system. We demonstrate that this problem can be formulated as a constrained convex optimization problem, and we use this formulation to study optimal harvesting and efficiency in the bacteriorhodopsin proton pump. We also solve the convex optimization problem in closed-form for three limiting regimes, as illustrated on two unicyclic models discussed above.
A key step in our analysis is to separate the dynamics of the system into separate contributions from fixed baseline processes and optimizable control processes. We note that, in stochastic thermodynamics, the baseline/control separation has been previously used to study autonomous Maxwellian demons [46, 47], counterdiabatic driving [48], and the cost of maintaining a nonequilibrium steady state [49, 50].
We derive a simplified bound on the maximum harvesting rate in (13), which is achieved in the limit of fast dissipation-less control. Interestingly, this expression involves a tradeoff between two terms, one information-theoretic and one thermodynamic. At first glance, this resembles information/free-energy tradeoffs characteristic of Maxwellian demons and other “information engines” [51, 52, 53, 54, 55, 56, 57, 58]. However, there are important differences. In a typical information engine, there is no external source of driving and information serves as fuel, which can be converted into ${\beta}^{-1}\ln 2$ of thermodynamic free energy per bit. In our case, there is an external source of free energy that in some cases can be harvested more effectively by reducing the system’s statistical entropy, e.g. by concentrating it on favorable states. Here, a bit of information can increase the harvesting rate by a very large amount (much larger than ${\beta}^{-1}\ln 2/\text{bit}$), and information acts more like a catalyst than a fuel [59, 60]. Loosely speaking, this is similar to how information encoded in the sequence of a metabolic enzyme is not itself fuel, but rather allows metabolism to harvest a large amount of fuel from elsewhere.
We finish by mentioning some connections to previous work and future directions. First, our approach may be related to prior work on optimizing power output and free energy transduction in steady-state molecular machines [61, 62, 63, 64, 65, 29, 66]. Here we consider the general problem of optimizing a set of control processes, given a fixed baseline and possible additional constraints on kinetics, topology, and thermodynamics. Previous work does not make the baseline/control distinction; instead, it is mostly concerned with the problem of optimizing system performance with respect to a small set of specific parameters or observables of interest, such as the location of free energy barriers [66, 62, 63, 65], efficiency [66], and the size of fluctuations [61, 64].
There is also an interesting relation between our work and flux balance analysis (FBA) [67, 68, 69]. The goal of FBA is to identify deterministic fluxes in biological metabolic systems that optimize biomass production, or other similar metrics of performance. This can be formulated as a linear program, which may include linear constraints that enforce thermodynamically favored reaction directions [69] (interestingly, in Ref. [70], the authors propose a version of FBA that also accounts for the entropy production rate). Our setting and optimization are different from FBA and its variants. We seek to optimize free energy harvesting at the stochastic level, and our objective involves nonlinear information-theoretic contributions to free energy. In addition, our optimization involves both the steady-state distribution $\bm{p}$ and fluxes ${\boldsymbol{J}}$, which allows us to optimize harvesting due to to internal transitions within coarse-grained mesostates, as in Fig. 4 (b). Nonetheless, investigating the relationship between our approach and FBA is an interesting direction for future work.
Another interesting direction for future work is to consider stochastic fluctuations of free energy harvesting. In particular, the thermodynamic uncertainty relation may be used to study tradeoffs between the entropy production rate, the average harvesting rate (the quantity $\dot{\mathcal{G}}^{\mathrm{tot}}$ considered here), and the fluctuations in the amount of harvested free energy [71, 64]. For biomolecular systems, large fluctuations in harvesting can lead to starvation, suggesting that minimizing fluctuations may be of significant biological importance.
Finally, an interesting direction for future work is to consider free energy harvesting in a system embedded in a fluctuating environment. For example, one may imagine a harvesting system in an environment with fluctuating sugar sources, or with a fluctuating amount of available light. In this setting, it is natural to optimize the harvesting rate under the topological constraint that control fluxes cannot directly change the state of the environment, for instance using bipartite models of Markovian dynamics [72]. It would be interesting to investigate how, under the optimal harvesting strategy, the information flow from the environment to the system varies with the abundance of free energy and complexity of the environment.